# Stabilization of active tissue deformation by a dynamic morphogen gradient

### v1

@@ Abstract
A key process during animal morphogenesis is oriented tissue deformation, which is often driven by internally generated active stresses. Yet, such active oriented materials are prone to well-known instabilities, raising the question of how oriented tissue deformation can be robust during morphogenesis. In a simple scenario, we recently showed that active oriented deformation can be stabilized by the boundary-imposed gradient of a scalar field, which represents, e.g., a morphogen gradient in a developing embryo. Here, we discuss a more realistic scenario, where the morphogen is produced by a localized source region, diffuses across the tissue, and degrades. Consistent with our earlier results, we find that oriented tissue deformation is stable in the gradient-extensile case, i.e. when active stresses act to extend the tissue along the direction of the gradient, but it is unstable in the gradient-contractile case. In addition, we now show that gradient-contractile tissues can not be stabilized even by morphogen diffusion. Finally, we point out the existence of an additional instability, which results from the interplay of tissue shear and morphogen diffusion. Our theoretical results explain the lack of gradient-contractile tissues in the biological literature, suggesting that the active matter instability acts as an evolutionary selection criterion.
@@ I Introduction
In animals, oriented tissue deformation can be affected by tissue-scale concentration patterns of proteins, which are often called morphogens. Such morphogen patterns are known to coordinate developmental processes on the tissue or even organism scale [1]. In many cases, they provide positional information [27, 3, 28]. Yet, in some cases they also provide orientational information that defines the axis of active stresses driving oriented tissue deformation [5, 29, 6, 12, 7], typically by guiding cellular polarity and cell rearrangements [2, 6, 12, 4, 30]. Such alignment of orientational information with morphogen gradients has recently also been discussed theoretically [31].
@@
In a recent publication [32], we theoretically showed that oriented tissue deformation can be stabilized by morphogen gradients. We used a hydrodynamic model for an active polar material and showed that the gradient of a scalar field, which describes the morphogen gradient, can stabilize the otherwise unstable material. If the scalar field advects with the material flows, we found that the stability of oriented material deformation depends on the coupling between scalar field gradient direction and deformation axis. In the gradient-extensile case, i.e. when anisotropic active stresses act to extend the material along the gradient direction, the scalar field can stabilize the active material deformation. However, in the gradient-contractile case, i.e. when anisotropic active stresses act to contract the material along the gradient direction, the deformation is unstable. However, one could always stabilize even a gradient-contractile system by sufficiently increasing morphogen diffusion.
@@
In our past publication [32], we have created a uniform scalar field gradient in a simple way through appropriate boundary conditions. While providing advantages for analytical calculations, this is quite different from the real mechanism creating morphogen gradients in developing biological tissues [33, 34, 1, 35, 36]. Morphogens do indeed advect with tissue flows [37], which we already considered in our previous work [32]. However, they are typically produced by some source region, diffuse across the tissue, and degrade [38, 36]. As a consequence, depending on the source region, this can lead to approximately exponentially decaying [33, 39, 40, 35] or periodic [41, 42] concentration profiles. However, it is not known how such morphogen dynamics would affect oriented deformation of an active material.
@@
In this article, we study the stability of active material deformation when accounting for more realistic morphogen dynamics (Figure 1A). We consider an active viscous material with local flow velocity $\bm{{\mathrm{v}}}(\bm{r},t)$, where active stresses are defined by the gradient of a morphogen field $c(\bm{r},t)$. We extend our previous work by more realistic morphogen dynamics. Specifically, the morphogen field $c(\bm{r},t)$ is secreted by a source region that is defined by a field $s(\bm{r},t)$, which is conserved and only advects with tissue flows. The morphogen field $c(\bm{r},t)$ not only advects with tissue flows, but also diffuses and degrades at constant rates everywhere.
@@
Using linear stability analysis, we find that also when including more realistic morphogen dynamics, gradient-extensile materials are stable, while gradient-contractile materials are unstable. Yet additionally, we also show that in contrast to a boundary-provided morphogen gradient [32], in the case of a secreted morphogen gradient, a gradient-contractile system can never be stabilized by diffusion. Finally, we find an additional instability in deforming systems for both gradient-extensile and gradient-contractile cases, which arises from the interaction between shear and diffusion, and which is unrelated to the Simha-Ramaswamy active matter instability.
@@
In the following, we first introduce the model in section II. We then study the case of a uniform gradient in both $s(\bm{r},t)$ and $c(\bm{r},t)$ for fixed system size in section III, for which the linear stability can be computed analytically. Next, in section IV, still for fixed system size, we study the more realistic case of a localized source $s(\bm{r},t)$, which gives rise to a non-uniform morphogen gradient $c(\bm{r},t)$. We then discuss the stability of freely deforming systems in section V. Finally, in section VI, we discuss our findings in the context of the biological literature.
@@
Figure 1: (A) Hydrodynamic model. A spatially varying source field $s(\bm{r},t)$ secretes a morphogen with concentration $c(\bm{r},t)$, which also diffuses and degrades. The gradient of $c$ creates extensile/contractile active anisotropic stresses, which drive viscous flows with velocity $\bm{{\mathrm{v}}}(\bm{r},t)$. These flows affect both source and morphogen fields through advection. (B) Typical orders of magnitude for tissue deformation rates (top), morphogen diffusion rates (middle), and morphogen degradation rates (bottom) from the biological literature. Blue marks are data from the fruit fly Drosophila melanogaster, and orange marks are data from zebrafish, Danio rerio. Tissue deformation rates in the fruit fly were estimated from: egg chamber [43, 44], dorsal thorax [12], pupal wing [8], hindgut [45, 5], renal tubules [7], and germ-band [11]; and in zebrafish from mesodermal explants [46] and the tailbud [47]. Morphogens: Decapentaplegic in the fruit fly [48, 33, 36]; Cyclops, Squint, Lefty1 and Lefty2 in zebrafish [34, 39].
@@ II.1 Hydrodynamic fields and bulk dynamics
We describe the tissue as a 2D active material, where we consider three interacting hydrodynamic fields: a scalar field $s(\bm{r},t)$, which describes the source, i.e. cells in the tissue which secrete the morphogen; a scalar field $c(\bm{r},t)$, which describes the morphogen concentration; and a vector field $\bm{{\mathrm{v}}}(\bm{r},t)$, which describes the tissue flows (Figure 1A). We find that literature values for the respective rates of tissue deformation, morphogen diffusion, and morphogen degradation are of roughly similar orders of magnitude (Figure 1B), suggesting that none of them can be easily neglected, and we thus include all three effects in our model.
The source field $s(\bm{r},t)$, which describes morphogen production, only advects with material flows:
@@
The morphogen field $c(\bm{r},t)$ follows a generalized conservation equation:
@@
Here, we have introduced production rate $k_{\mathrm{p}}$, degradation rate $k_{\mathrm{d}}$, and diffusion coefficient $D$.
@@
This expression for the anisotropic active stress is the same as in Active Model H [49, 50]. Its direction is determined by the sign of the coefficient $\alpha$: For $\alpha<0$, the activity is gradient-extensile, i.e. it acts to extend the material along the local $c$ gradient direction and contract perpendicularly; while for $\alpha>0$, the activity is gradient-contractile, i.e. it acts to contract the material along the $c$ gradient direction and extend perpendicularly (Figure 1A).
The expression for the active stress, Eq. (6), is motivated by what is known about developing biological tissues. Specifically, anisotropic internal stresses are usually generated by an anisotropic distribution of cytoskeletal elements within cells. Such an anisotropic distribution is created by cell polarity, which in turn can be controlled by morphogen gradients [12, 51, 52, 7, 30]. Here, we assume that the time scales involved in these different couplings are fast enough, so that we directly couple the generated active stresses to the morphogen gradient. Eq. (6) represents the lowest-order term for such a coupling [49].
@@ II.2 Boundary Conditions
    $\frac{1}{L_{x}}\,\frac{\mathrm{d}L_{x}}{\mathrm{d}t}=-\frac{\alpha}{4\eta}\int{\big{[}(\partial_{x}c)^{2}-(\partial_{y}c)^{2}\big{]}\;\mathrm{d}A},$ (7)
@@
We investigate the stability of two stationary states: (i) a state where a linear source profile $s$ gives rise to a linear morphogen profile $c$, and (ii) a state where a localized source $s$ generates a spatially decaying morphogen field $c$. In both cases, the amplitude of the source field is defined by a constant $s_{b}$. In case (i), which we study in section III, the boundary conditions are periodic with the exception of the vertical boundaries for the scalar fields $s$ and $c$, where we set for all $y\in[0,L_{y})$:
@@ II.3 Dimensionless equations
  • We choose the unit of length such that the system area is one $L_{x}L_{y}=1$.
  • We choose the unit of the morphogen field $c$ such that $k_{\mathrm{p}}=k_{\mathrm{d}}$.
@@
The last condition allows to compare the linear stability of systems with different source profiles, beacuse the reference time scale is always the respective free deformation rate.
@@
00 0 $=\frac{1}{4}\partial_{i}^{2}{\mathrm{v}}_{j}-\partial_{j}\Pi^{\prime}+\mathrm{sgn}(\alpha)\,\partial_{i}\big{[}(\partial_{i}c)(\partial_{j}c)\big{]},$ (12)
@@
Here, we have introduced the sign function, $\mathrm{sgn}(\alpha):=\alpha/|\alpha|$, and have set $\Pi^{\prime}:=\Pi+\mathrm{sgn}(\alpha)(\partial_{k}c)^{2}/2$. The offset introduced at the boundary conditions in Eqs. (8) and (9) is rescaled to $1$ for both scalar fields.
@@ III Linear source profile
Figure 2: Stability of a fixed-size system with a linear source profile: gradient-contractile systems (orange) are unstable, whereas gradient-extensile systems (green) are marginally stable. (A) The linear source profile, $s_{0}(x)=x$ in dimensionless units. In the stationary state, $c_{0}(x)=s_{0}(x)$. (B) The perturbation growth rates $\omega$ depending on wave vector angles $\phi$. There are 2 branches for each of the 4 cases reported (see legend). (C) Effect of morphogen diffusion rate $Dq_{0}^{2}$ (with $q_{0}=2\pi$) on the maximal perturbation growth rates $\omega_{\mathrm{max}}$ of the linear systems. Inset: same data in log-log scaling for the gradient-contractile case.
We first discuss the limit without diffusion, $D=0$. When we subtract the linearized versions of the source dynamics, Eq. (10), from the linearized version of the morphogen dynamics, (11), we obtain:
@@
Thus, any difference between morphogen field $c$ and source field $s$ decays at rate $k_{\mathrm{d}}$ (dashed horizontal lines in Figure 2B). In addition, the combination of morphogen dynamics and generated active flows, Eqs. (11)–(13), leads to the same results as for a single scalar field discussed in earlier publications [50, 32] (Appendix D.1). This means that the perturbation growth rate $\omega_{\mathrm{act}}$ due to the active flows depends on the angle of the Fourier mode $\phi$ as $\omega_{\mathrm{act}}:=4\,\mathrm{sgn}(\alpha)\sin^{2}{\phi}$, as indicated by the dashed curves in Figure 2B. Thus, without diffusion, the system is marginally stable in the gradient-extensile case, where the maximal perturbation growth rate is zero (green dashed curve), and unstable in the gradient-contractile case, where the maximum perturbation growth rate is strictly positive (orange dashed curve).
Even in the presence of diffusion, $D>0$, a gradient-contractile system remains always unstable, although the maximal perturbation growth rate decreases with diffusion (orange curve in Figure 2C & inset). This is in contrast to earlier findings for a system with a single scalar field $c$ without source, where we found that diffusion can stabilize the system in the gradient-contractile case. To understand where the difference comes from, we consider the limit of very large $D$. The linearized version of the morphogen dynamics, Eq. (11), in Fourier space reads:
@@
The flow field $\bm{\delta{\mathrm{v}}}$ enters on the left-hand side due to advection, where $\partial_{x}c_{0}=1$ in dimensionless units. The flow $\bm{\delta{\mathrm{v}}}$ is actively created by the morphogen field itself through Eqs. (12) and (13), and their $x$ component in Fourier space reads: $\delta{\mathrm{v}}_{x}=-\omega_{\mathrm{act}}\delta c$, with the active flow time scale $\omega_{\mathrm{act}}$ introduced before (Appendix D). Denoting the minimal non-zero wave vector magnitude in our system by $q_{\mathrm{min}}$, we take the limit where $Dq_{\mathrm{min}}^{2}$ is much larger than all other relevant time scales in the system. In this case, according to Eq. (17), the Fourier amplitudes $\delta c$ for each wave vector relax adiabatically fast towards
@@
This effectively corresponds to a system where we again just have a single scalar field $s$, which advects with active tissue flows $\bm{{\mathrm{v}}}$, but does not diffuse. The active tissue flows are generated by the perturbations of $c$, and their $x$ component is given by $\delta{\mathrm{v}}_{x}=-\omega_{\mathrm{act}}\delta c\simeq-(\omega_{\mathrm{act}}k_{\mathrm{d}}/Dq^{2})\delta s$. Inserting this in the linearized version of the source dynamics in Fourier space, we get:
$\frac{\partial\delta s}{\partial t}=-\delta{\mathrm{v}}_{x}=\frac{\omega_{\mathrm{act}}k_{\mathrm{d}}}{Dq^{2}}\delta s=\frac{4k_{\mathrm{d}}\mathrm{sgn}(\alpha)\sin^{2}{\phi}}{Dq^{2}}\delta s.$ (19)
Thus, the sign of the growth rate of $\delta s$ just depends on the sign of $\alpha$. As a consequence, the system is unstable for a gradient-contractile system ($\alpha>0$), even in the limit of a large diffusion constant. The reason for this is that the source field itself advects with the active flows, but does not diffuse. For the same reason, any gradient-extensile system remains only marginally stable even in the presence of morphogen diffusion.

### v2

# Stabilization of active tissue deformation by a dynamic signaling gradient

@@ Abstract
A key process during animal morphogenesis is oriented tissue deformation, which is often driven by internally generated active stresses. Yet, such active oriented materials are prone to well-known instabilities, raising the question of how oriented tissue deformation can be robust during morphogenesis. Here we study under which conditions active oriented deformation can be stabilized by the concentration pattern of a signaling molecule, which is secreted by a localized source region, diffuses across the tissue, and degrades. Consistent with earlier results, we find that oriented tissue deformation is always unstable in the gradient-contractile case, i.e. when active stresses act to contract the tissue along the direction of the signaling gradient, and we now show that this is true even in the limit of large diffusion. However, active deformation can be stabilized in the gradient-extensile case, i.e. when active stresses act to extend the tissue along the direction of the signaling gradient. Specifically, we show that gradient-extensile systems can be stable when the tissue is already elongated in the direction of the gradient. We moreover point out the existence of a formerly unknown, additional instability of the tissue shape change. This instability results from the interplay of active tissue shear and signal diffusion, and it indicates that some additional feedback mechanism may be required to control the target tissue shape. Taken together, our theoretical results provide quantitative criteria for robust active tissue deformation, and explain the lack of gradient-contractile systems in the biological literature, suggesting that the active matter instability acts as an evolutionary selection criterion.
@@ I Introduction
In animals, oriented tissue deformation can be affected by tissue-scale concentration patterns of signaling molecules. Such patterns of signaling molecules are known to coordinate developmental processes on the tissue or even organism scale [1]. Signaling molecules can provide positional information, in which case they are called morphogens [27, 3, 28, 29, 30, 31]. Yet, in some cases, concentration gradients of signaling molecules also provide orientational information, which defines the axis of active stresses driving oriented tissue deformation [5, 32, 6, 12, 7], typically by guiding cellular polarity and cell rearrangements [2, 6, 12, 4, 33]. Such alignment of orientational information with signaling gradients has recently also been discussed theoretically [34].
@@
In a recent publication [35], we theoretically showed that oriented tissue deformation can be stabilized by signaling gradients. We used a hydrodynamic model for an active polar material, and showed that the gradient of a scalar signaling field can stabilize the otherwise unstable material. We showed that if the scalar field advects with the material flows [36, 37], the stability of oriented material deformation depends on the coupling between scalar field gradient direction and deformation axis. In the gradient-extensile case, i.e. when anisotropic active stresses act to extend the material along the gradient direction, the scalar field can stabilize the active material deformation. However, in the gradient-contractile case, i.e. when anisotropic active stresses act to contract the material along the gradient direction, the deformation is unstable. Yet, one could always stabilize even a gradient-contractile system by sufficiently increasing the diffusion of the signaling molecule.
@@
In our past publication [35], we have created a uniform scalar field gradient in a simple way through appropriate boundary conditions. While providing advantages for analytical calculations, this is quite different from the real mechanism creating signaling gradients in developing biological tissues [38, 28, 29, 1, 39, 40]. Signaling molecules are typically produced by some source region, diffuse across the tissue, and degrade [41, 40]. As a consequence, depending on the source region, this can lead to approximately exponentially decaying [38, 42, 43, 39] or periodic [44, 45] concentration profiles. However, it is not known how such signaling molecule dynamics would affect stability of active oriented tissue deformation.
@@
In this article, we study the stability of active tissue deformation when accounting for more realistic signaling dynamics (Figure 1A). We consider an active viscous tissue with local flow velocity $\bm{{\mathrm{v}}}(\bm{r},t)$, where active stresses are defined by the gradient of a signaling molecule concentration field $c(\bm{r},t)$. The signaling field $c(\bm{r},t)$ is in turn secreted by a source region that is defined by a field $s(\bm{r},t)$, which is conserved and only advects with tissue flows. The signaling field $c(\bm{r},t)$ not only advects with tissue flows, but also diffuses and degrades at constant rates everywhere.
@@
Using linear stability analysis, we find that gradient-contractile tissues are always unstable, consistent with our earlier findings [35]. Yet, we also show that for our more realistic signaling dynamics, a gradient-contractile system can never be stabilized by diffusion. We further show that gradient-extensile systems can be stable, in particular if the tissue is longer in the direction of the gradient than transversal to it. Finally, as soon as the system deforms, we find that an additional instability can appear for both gradient-extensile and gradient-contractile cases, arising from the interaction between shear and diffusion. This instability may point to the necessity of an additional control mechanism for the final tissue dimensions.
@@
In the following, we first introduce the model in section II. We then study the case of a uniform gradient in both $s(\bm{r},t)$ and $c(\bm{r},t)$ for fixed system size in section III, for which the linear stability can be computed analytically. Next, in section IV, still for fixed system size, we study the more realistic case of a localized source $s(\bm{r},t)$, which gives rise to a non-uniform signaling gradient $c(\bm{r},t)$. We then discuss the stability of freely deforming systems in section V. Finally, in section VI, we discuss our findings in the context of the biological literature.
@@
Figure 1: (A) Hydrodynamic model. A spatially varying source field $s(\bm{r},t)$ secretes a signaling molecule with concentration $c(\bm{r},t)$, which also diffuses and degrades. The gradient of $c$ creates extensile/contractile active anisotropic stresses, which drive viscous flows with velocity $\bm{{\mathrm{v}}}(\bm{r},t)$. These flows affect both source and signaling fields through advection. (B) Typical orders of magnitude for tissue deformation rates $\tilde{v}^{0}_{xx}$ (top), signal diffusion rates $D/L^{2}$, where $D$ is the diffusion coefficient and $L$ is the linear dimension of the corresponding tissue (middle), and signal degradation rates $k_{\mathrm{d}}$ (bottom) from the biological literature. Blue marks are data from the fruit fly Drosophila melanogaster, and orange marks are data from zebrafish, Danio rerio. Tissue deformation rates in the fruit fly were estimated from image time series of: egg chamber [46, 47], dorsal thorax [12], pupal wing [8], hindgut [48, 5], renal tubules [7], and germ-band [11]; and in zebrafish from mesodermal explants [49] and the tailbud [50]. Signaling molecules: Decapentaplegic in the fruit fly [51, 38, 40]; Cyclops, Squint, Lefty1 and Lefty2 in zebrafish [29, 42].
@@ II.1 Hydrodynamic fields and bulk dynamics
We describe the tissue as a 2D active material, where we consider three interacting hydrodynamic fields: a scalar field $s(\bm{r},t)$, which describes the source, i.e. cells in the tissue which secrete the signaling molecule; a scalar field $c(\bm{r},t)$, which describes the signaling molecule concentration; and a vector field $\bm{{\mathrm{v}}}(\bm{r},t)$, which describes the tissue flows (Figure 1A). We find that literature values for the respective rates of tissue deformation $\tilde{v}^{0}_{xx}$, signaling molecule diffusion $D/L^{2}$, where $D$ is the diffusion coefficient and $L$ denotes the linear dimension of the tissue, and signaling molecule degradation $k_{\mathrm{d}}$ are of roughly similar orders of magnitude (Figure 1B), suggesting that none of them can be easily neglected, and we thus include all three effects in our model.
The source field $s(\bm{r},t)$, which describes signaling molecule secretion, only advects with material flows:
@@
The signaling field $c(\bm{r},t)$ follows a generalized conservation equation:
@@
Here, we have introduced production rate $k_{\mathrm{p}}$, degradation rate $k_{\mathrm{d}}$, and diffusion coefficient $D$. Note that there are several intra- and extra-cellular mechanisms of signaling molecule transport [29] – here we describe the emergent effective transport on large length scales by advection and diffusion.
@@
This expression for the active anisotropic stress is the same as in Active Model H [52, 53]. Its direction is determined by the sign of the coefficient $\alpha$: For $\alpha<0$, the activity is gradient-extensile, i.e. it acts to extend the material along the local $c$ gradient direction and contract perpendicularly; while for $\alpha>0$, the activity is gradient-contractile, i.e. it acts to contract the material along the $c$ gradient direction and extend perpendicularly (Figure 1A).
The expression for the active stress, Eq. (6), is motivated by what is known about developing biological tissues. Specifically, anisotropic internal stresses are usually generated by an anisotropic distribution of cytoskeletal elements within cells. Such an anisotropic distribution is created by cell polarity, which in turn can be controlled by signaling gradients [12, 54, 55, 7, 33]. Rather than explicitly including cell polarity as polar field [35], we assume here for simplicity a direct coupling of active stresses to the signaling gradient. This corresponds to the limit where the time scales involved in couplings active stresses to the signaling gradient are fast as compared to the other time scales of the system. While this is clearly a simplification, we do not expect the fundamental results of the analysis do not differ much in this limit, as suggested by our earlier work [35]. Note that Eq. (6) represents the lowest-order term for a coupling of an active, anisotropic stress to a scalar field [52].
@@ II.2 Boundary Conditions
    $\frac{1}{L_{x}}\,\frac{\mathrm{d}L_{x}}{\mathrm{d}t}=-\frac{\alpha}{4\eta L_{x}L_{y}}\int{\big{[}(\partial_{x}c)^{2}-(\partial_{y}c)^{2}\big{]}\;\mathrm{d}A},$ (7)
@@
We investigate the stability of two stationary states: (i) a state where a linear source profile $s$ gives rise to a linear signaling profile $c$, and (ii) a state where a localized source $s$ generates a spatially decaying signaling field $c$. In both cases, the amplitude of the source field is defined by a constant $s_{b}$. In case (i), which we study in section III, the boundary conditions are periodic with the exception of the vertical boundaries for the scalar fields $s$ and $c$, where we set for all $y\in[0,L_{y})$:
@@ II.3 Dimensionless equations
  • We choose the unit of length such that the system area is one, $L_{x}L_{y}=1$.
  • We choose the unit of the signaling field $c$ such that $k_{\mathrm{p}}=k_{\mathrm{d}}$.
@@
The last condition allows to compare the linear stability of systems with different source profiles, because the reference time scale is always the respective free deformation rate.
@@
$0$ $=\frac{1}{4}\partial_{i}^{2}{\mathrm{v}}_{j}-\partial_{j}\Pi^{\prime}+\mathrm{sgn}(\alpha)\,\partial_{i}\big{[}(\partial_{i}c)(\partial_{j}c)\big{]},$ (12)
@@
Here, we have introduced the sign function, $\mathrm{sgn}(\alpha):=\alpha/|\alpha|$, and have set $\Pi^{\prime}:=\Pi+\mathrm{sgn}(\alpha)(\partial_{k}c)^{2}/2$. The offset introduced at the boundary conditions in Eqs. (8) and (9) is rescaled to 11 for both scalar fields.
@@ III Linear source profile
[…]
We first discuss the limit without diffusion, $D=0$. When we subtract the linearized versions (Appendix C) of source and signaling dynamics, Eqs. (10) and (11), we obtain:
@@
Thus, any difference between signaling field $c$ and source field $s$ decays at rate $k_{\mathrm{d}}$. In addition, the combination of signaling dynamics and generated active flows, Eqs. (11)–(13), leads to the same results as for a single scalar field discussed in earlier publications [53, 35] (Appendix D.1). Specifically, for the combined perturbation field $\delta c_{\mathrm{eff}}:=\delta c+k_{\mathrm{d}}/\omega_{\mathrm{act}}\delta s$, we obtain (Eq. (72) in Appendix D):
$\frac{\partial\delta c_{\mathrm{eff}}}{\partial t}=\omega_{\mathrm{act}}\,\delta c_{\mathrm{eff}},$ (17)
where $\omega_{\mathrm{act}}:=4\,\mathrm{sgn}(\alpha)\sin^{2}{\phi}$ is the perturbation growth rate created by the active flows. It takes the same form without source field [35]. Thus, without diffusion, the system is marginally stable in the gradient-extensile case, i.e. the maximal perturbation growth rate is zero, and unstable in the gradient-contractile case.
In the presence of diffusion, $D>0$, gradient-extensile systems remain marginally stable (Figure 2C). However, gradient-contractile systems can not be stabilized by diffusion, even though the maximal perturbation growth rate decreases with diffusion (Figure 2B,D). This is in contrast to earlier findings for a system with a boundary-imposed signaling gradient, where we found that diffusion can stabilize gradient-contractile systems [35]. To understand where the difference comes from, we consider the limit of very large $D$. The linearized version of the signaling dynamics, Eq. (11), in Fourier space reads:
@@
The flow field $\bm{\delta{\mathrm{v}}}$ enters on the left-hand side due to advection, where $\partial_{x}c_{0}=1$ in dimensionless units. The flow $\bm{\delta{\mathrm{v}}}$ is actively created by the signaling field itself through Eqs. (12) and (13), and its $x$ component in Fourier space reads: $\delta{\mathrm{v}}_{x}=-\omega_{\mathrm{act}}\delta c$, with the active flow time scale $\omega_{\mathrm{act}}$ introduced before. Denoting the minimal non-zero wave vector magnitude in our system by $q_{\mathrm{min}}$, we take the limit where $Dq_{\mathrm{min}}^{2}$ is much larger than all other relevant time scales in the system. In this case, according to Eq. (18), the Fourier amplitude $\delta c$ for each wave vector relaxes adiabatically fast towards
@@
This effectively corresponds to a system where we again just have a single scalar field $s$, which advects with active tissue flows $\bm{{\mathrm{v}}}$, but does not diffuse. The active tissue flows are generated by the perturbations of $c$, and their $x$ component is given by $\delta{\mathrm{v}}_{x}=-\omega_{\mathrm{act}}\delta c\simeq-(\omega_{\mathrm{act}}k_{\mathrm{d}}/Dq^{2})\delta s$. Inserting this into the linearized version of the source dynamics in Fourier space, we obtain (compare to Eq. (17)):
$\frac{\partial\delta s}{\partial t}=-\delta{\mathrm{v}}_{x}=\frac{\omega_{\mathrm{act}}k_{\mathrm{d}}}{Dq^{2}}\delta s.$ (20)
Thus, the perturbation growth rate is given by the growth rate without diffusion, $\omega_{\mathrm{act}}$, multiplied by $k_{\mathrm{d}}/Dq^{2}$ (Figure 2B). As a consequence, gradient-contractile systems are unstable even in the limit of a large diffusion constant. The reason for this is that the source field itself advects with the active flows, but does not diffuse. For the same reason, gradient-extensile systems remains only marginally stable even in the presence of signaling molecule diffusion.